OpenStudy (anonymous):

I want to know is my answer right? some check it plz! Solve using the substitution method. 1/2x-1/5y = 7, 1/6x+ 2/3= -4 my answer Is this right? 1/2x-1/5y = 7, ---1) 1/6x+ 2/3= -4 ---2) Rearrange equation 2) 1/6x+ 2/3= -4 1/6x = -4 -2/3 x = -28 substitute x = -28 into 1) 1/2 x -28 -1/5y =7 -14 - 1/5y = 7 1/5y = -14 -7 y = 105

OpenStudy (anonymous):

1/2x-1/5y = 7, ---1) 1/6x+ 2/3= -4 ---2) Rearrange equation 2) 1/6x+ 2/3= -4 1/6x = -4 -2/3 x = -28 substitute x = -28 into 1) 1/2 x -28 -1/5y =7 -14 - 1/5y = 7 1/5y = -14 -7 y = 105

OpenStudy (anonymous):

is this right??

hero (hero):

Wait...the second equation, doesn't have a y in it...Is that the way it's supposed to be?

OpenStudy (anonymous):

1/2x-1/5y = 7, 1/6x+ 2/3= -4

OpenStudy (anonymous):

like this

hero (hero):

So, the second equation has no y in it, correct?

OpenStudy (anonymous):

yes

hero (hero):

\[\frac {x}{2}-\frac{y}{5}=7\] and \[\frac{x}{6}-\frac{2}{3}=-4\]

hero (hero):

In the first equation multiply both sides by 10: 5x-2y = 70 In the second equation, multiply both sides by 6: x+4 = -24

OpenStudy (anonymous):

What you wrote: -14 - 1/5y = 7 1/5y = -14 -7 y = 105 What I think would be correct: -14 - 1/5y = 7 -1/5y=7-14 -1/5y=-7 multiply -5 to each side y=35 Then you have to find x.

hero (hero):

I messed that all up

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

(It was a little bit jumbled so it took me quite a while to follow what you did, but everything was fine until the end.)

OpenStudy (anonymous):

Hahahha, um, do you get what I did?

OpenStudy (anonymous):

no!

hero (hero):

That's what I get

hero (hero):

Wow...that's approximately 35

hero (hero):

so maybe nitaleung's methods are accurate

OpenStudy (anonymous):

LOL okay I'm confusing myself too. I should have added 14 instead of subtract... -14 - 1/5y = 7 -1/5y=7+14 -1/5y=21 multiply -5 to each side y=105 Oh pft, I'm such a doofus. --But I wonder how Hero got his answer then?? Donnice, you're straight on correct, ignore me please.

hero (hero):

I'll recalculate....

hero (hero):

y = 105

hero (hero):

I did a wrong step somewhere the first time

OpenStudy (anonymous):

:) great

OpenStudy (anonymous):

so my answer right

OpenStudy (anonymous):

U may always check this sort of thing by feeding your equations to the Wolfram god.... http://www.wolframalpha.com/input/?i=x%2F2+-+y%2F5+%3D+7%2C+x%2F6%2B+2%2F3%3D+-4